OpenStudy (acoleman3):

Find the magnitude & direction of the vector JK given points J(1,0) & K(-2,3). Select one: a. 32 √ 32 ; 135 degrees b. 43 √ 43 ; 140 degrees c. 6 ; 135 degrees d. 55 √ 55 ; 150 degrees

OpenStudy (acespeedfighter):

@Moderator

OpenStudy (acespeedfighter):

@OpenStudy

OpenStudy (acespeedfighter):

@Admin

OpenStudy (sshayer):

vector JK=(-2-1,3-0)=(-3,3) \[magnitude=\left| JK \right|=\sqrt{\left( -3 \right)^2+3^2}=\sqrt{18}=3\sqrt{2}\]

OpenStudy (sshayer):



OpenStudy (sshayer):

correction \[\tan \theta=\frac{ 3 }{ -3 }=-1\]